# Graph Network Simulator MGN MeshGraphNet

### v1

@@ 4 Experiments
Figure 5: Left: Rollout error measured by the RMSE of the predicted nodes positions. ROBIN surpasses the accuracy of the baselines on all three datasets BendingBeamImpactPlate and  Deformingplate. Right: Comparison of inference time and error of ROBIN and its variants on BendingBeam. ROBIN achieves the same accuracy as conventional diffusion inference (ROBIN - $m20$), while the inference speed is close the one step variant (ROBIN - OS). Reducing the denoising step $K$, denoted as ROBIN - $K10$ and ROBIN - $K5$, trades accuracy for speed.
@@
Datasets. We evaluate our model on the three different datasets, namely BendingBeam, ImpactPlate [22] and DeformingPlate [5]. We introduce the BendingBeam dataset (Figure 4a)), featuring quasi-static, geometrically non-linear deformations of beams with high aspect ratios. The setup challenges models to capture global deformations via broad receptive fields and resolve high spatial frequencies due to locally thin, low-stiffness regions. In ImpactPlate (Figure 4 b)), the models must learn flexible dynamics with varying material parameters and accurately resolve very localized deformation at the contact point. DeformingPlate (Figure 4 c)) considers quasi-static contact simulations induced by scripted actuators that deform $3$D plates consisting of nonlinear, hyperelastic material.
@@
Experimental setup. For all tasks, we target the displacement residual of the node positions with respect to the next time step $\Delta\mathbf{u}^{t}_{i,0}=\mathbf{x}^{t+1}_{i}-\mathbf{x}^{t}_{i}$ during denoising. We additionally denoise the von Mises stress $\sigma^{t}_{\text{vMises},i,0}$ directly without residuals to gain further insight into the dynamics of the three experiments. ROBIN uses $K=20$ denoising steps and a denoising stride of $m=1$ by default. The task-specific features are specified in Appendix C, listed in Table 2. We measure the prediction error using an RMSE, as specified in Appendix C. Appendix C provides also information about additional settings of ROBIN, including training details and hyperparameters.
@@
Baselines. We compare our model with two prominent baselines for nonlinear deformation simulations, namely MGNs and Hierarchical Contact Mesh Transformes. In Appendix D, we describe the baselines and their setup in more detail.
Variants. We demonstrate that trained ROBINs can easily switch between different rollout modes by varying the denoising stride $m$. We compare ROBIN to conventional autoregressive diffusion inference (ROBIN - $m20$), i.e., $m=20$ denoising steps per time step, and to an intermediate variant (ROBIN - $m5$) with a denoising stride of $m=5$. Additionally, without retraining, we can run ROBIN as one step model, namely ROBIN - 20/OS, which uses the clean sample $\Delta\tilde{\mathbf{u}}^{t_{b}}_{0}\approx\Delta\tilde{\mathbf{u}}^{t_{b}}_{0|K}$ of the first denoising step as prediction. Finally, we train ROBIN with fewer diffusion steps $K=5$ (ROBIN - $K5$) and $K=10$ (ROBIN - $K10$), running both with a denoising stride of $m=1$.
@@
Ablations. We ablate key components of ROBIN to assess their impact. No diffusion trains the AMPN as a one-step autoregressive model with an MSE loss. No hierarchy disables hierarchical message passing, operating only on the fine mesh $G^{0}$. State prediction replaces residual-based prediction with direct denoising of $\mathbf{u}^{t}_{i,k}$ instead of $\Delta\mathbf{u}^{t}_{i,k}$. No shared layer uses $15$ unshared message passing layers. Appendix D provides additional implementation and training details.
@@ 5 Results
Baselines. Figure 5 a) comparse the rollout RMSE of ROBIN to HCMT and MGN. ROBIN yields substantial error reductions for the prominent ImpactPlate and DeformingPlate datasets, and achieves an even more remarkable improvement on BendingBeam. Figure 6 demonstrates that ROBIN is able to propagate local boundary conditions accross the part for accurate predicitons of the global part deformation. In addition, ROBIN resolves the non-linear bending curve of the FEM. This requires to combine geometric features at different scales, such as global part dimensions with local thin walls, for an accurate prediction of the global part stiffness.
@@
Figure 6: Comparison of the predicted rollout deformations and von Mises stresses (color, yellow is large) on BendingBeam between a) the FEM, b) ROBIN, c) HCMT, and d) MGN. ROBIN is able to accurately reproduce the FEM results. Both HCMT and MGN are unable to resolve global deformation modes, illustrating the importance of the AMPN for global message propagation.
@@
Inference speed. In Figure 5 b), the rollout RMSE is plotted over the inference time for the BendingBeam dataset. ROBIN is an order of magnitude faster than conventional autoregressive diffusion inference, denoted as ROBIN - $m20$, without compromising accuracy. Most notably, the fastet variant ROBIN - OS, which is a one step inference variant of ROBIN, significantly outperform the accuracy of the baselines (cf. Figure 5 a)). Despite $K=20$ denoising steps, ROBIN requires only $\approx 71\%$ longer than the one step variant, due to the parallel denoising scheme ROBI. By training ROBIN with fewer denoising steps, we obtain models, such as ROBIN - $K10$ and ROBIN - $K5$, whose speed and accuracy lies between the one-step variant and ROBIN. We obtain similar results for ImpactPlate and DeformingPlate, which is visualized in Figure 9 in Appendix E.
@@
Ablations. As visualized in Figure 7, the ablation results demonstrate the effectiveness of the individual components of ROBIN. Non-shared layers decrease the performance significantly on BendingBeam, due the the broader receptive field. The large error increase of the State prediction ablations indicates that residual prediction are crucial for Lagrangian simulations. ROBIN outperforms the non-hierarchical architecture variant as well as the non-diffusion variant on all datasets, demonstrating the synergy between DDPMs and AMPNs.
@@
Figure 7: Rollout RMSE of the displacement predictions. Shared layers are crucial for BendingBeam to increase the receptive field. Replacing residual predictions with state predictions, hierarchal architectures with non-hierarchical architectures, and diffusion with non-diffusion architectures significantly decrease the accuracy across all datasets.
@@ 6 Conclusion
Limitations and Future Work. ROBIN currently does not possess SO(3) equivariance, limiting its applicability to predictions involving tensor fields or anisotropic materials. We focus on DDPM, while other diffusion formulations or denoising schedules could provide valuable insights and further enhance performance. Similarly, our experiments cover nonlinear solid mechanics, and extensions to other domains, such as fluid dynamics, are a promising direction. Lastly, ROBIN’s combination of fast inference and high accuracy opens opportunities for accelerating multi-stage design and optimization workflows.
@@ Acknowledgements
This work is part of the DFG AI Resarch Unit 5339 regarding the combination of physics-based simulation with AI-based methodologies for the fast maturation of manufacturing processes. The financial support by German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) is gratefully acknowledged. The authors acknowledge support by the state of Baden-Württemberg through bwHPC, as well as the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the German Federal Ministry of Education and Research. This work is supported by the Helmholtz Association Initiative and Networking Fund on the HAICORE@KIT partition.

### v2

@@ 4 Experiments
Datasets. We evaluate our model on the three different datasets, namely BendingBeam, ImpactPlate [22] and DeformingPlate [5]. We introduce the BendingBeam dataset (Figure˜3a)), featuring quasi-static, geometrically non-linear deformations of beams with high aspect ratios. The setup challenges models to capture global deformations via broad receptive fields and resolve high spatial frequencies due to locally thin, low-stiffness regions. In ImpactPlate (Figure˜3 b)), the models must learn flexible dynamics with varying material parameters and accurately resolve very localized deformation at the contact point. DeformingPlate (Figure˜3 c)) considers quasi-static contact simulations induced by scripted actuators that deform 33D plates consisting of nonlinear, hyperelastic material.
@@
Experimental setup. For all tasks, we target the displacement residual of the node positions with respect to the next time step $\Delta\mathbf{u}^{t}_{i,0}=\mathbf{x}^{t+1}_{i}-\mathbf{x}^{t}_{i}$ during denoising. We additionally denoise the von Mises stress $\sigma^{t}_{\text{vMises},i,0}$ directly without residuals to gain further insight into the dynamics of the three experiments. ROBIN uses $K=20$ denoising steps and a denoising stride of $m=1$ by default. The task-specific features are specified in Appendix˜C, listed in Table˜2. We measure the prediction error using an RMSE, as specified in Appendix˜C. Appendix˜C also provides information about additional settings of ROBIN, including training details and hyperparameters.
@@
Baselines. We compare our model with three prominent baselines for nonlinear deformation simulations, namely MGNsHierarchical Contact Mesh Transformers, and Bi-Stride Multi-Scale GNNs. Detailed setups are provided in Appendix˜D.
[…]
@@
Ablations. We ablate key components of ROBIN to assess their impact. 10 diffusion steps and 5 diffusion steps train with reduced diffusion length $K$. W/o diffusion trains the AMPN as a one-step autoregressive model with MSE loss. W/o hierarchy disables hierarchical message passing, operating only on the fine mesh $G^{0}$. State prediction replaces residual-based prediction with direct denoising of $\mathbf{u}^{t}_{i,k}$ instead of $\Delta\mathbf{u}^{t}_{i,k}$. W/o shared layer uses $15$ unshared message passing layers. HCMT model replaces AMPNs by HCMTs as the hierarchical model for ROBIN. Appendix˜D provides additional implementation and training details.
Generalization to large meshes. To assess upscaling, we introduce BendingBeamLarge with meshes on average over ten times larger than in BendingBeam. We compare fine-tuning a pre-trained ROBIN to training from scratch, showing rapid transfer to substantially larger meshes without architectural changes, enabled by shared blocks in the AMPNs that accommodate varying hierarchy depths.
@@ 5 Results
Baselines. Figure˜4 a) compares the rollout RMSE of ROBIN against HCMT,  MGN, and BSMS. ROBIN yields substantial error reductions for the prominent ImpactPlate and DeformingPlate datasets, and yields an even larger improvement on BendingBeam. Figure˜5 demonstrates that ROBIN is able to propagate local boundary conditions across the part for accurate predictions of the global part deformation. In addition, ROBIN resolves the non-linear bending curve of the FEM. This requires fusing geometric features across scales, e.g., global part dimensions together with local thin walls, for an accurate prediction of the global part stiffness. All baselines struggle to reproduce the FEM results, particularly the global deformation modes.
Figure 4: Left: Rollout error measured by the RMSE of the predicted nodes positions. ROBIN surpasses the accuracy of the baselines HCMTMGN, and BSMS on all three datasets BendingBeamImpactPlate and  DeformingPlate. Right: Comparison of inference time and error of ROBIN and its variants on BendingBeam. The default variant of ROBIN (1/20) achieves the same accuracy as conventional diffusion inference (20/20), while the inference speed is close to the one step variant (1/1). Reducing the truncation step $k_{\text{tr}}$ trades accuracy for speed.
@@
Figure 5: Comparison of the predicted rollout deformations and von Mises stresses (color, yellow is high) on BendingBeam between a) the FEM, b) ROBIN, c) HCMT, d) MGN, and e) BSMS. ROBIN is able to accurately reproduce the FEM results. Neither HCMTMGN nor BSMS are able to resolve global deformation modes, illustrating the importance of the AMPN for global message propagation.
@@
Inference speed. Figure˜4 b) plots rollout RMSE versus inference time on BendingBeam. The default variant (1/20) of ROBIN is about an order of magnitude faster than conventional autoregressive diffusion inference (20/20), without compromising accuracy. Decreasing the truncation step $k_{\text{tr}}$ (i.e., fewer diffusion steps per physical step at inference) trades rollout accuracy for speed. Nevertheless, the fastest variant (1/1), which can be seen as a one-step model variant of ROBIN, still significantly outperforms the accuracy of the baselines (cf. Figure˜4 a)). Most notably, ROBIN (1/20) requires only $\approx 58\%$ more time than the one-step variant (1/1), despite performing K=20K{=}20 denoising steps per time step, highlighting the efficiency of the parallel denoising scheme ROBI. Similar trends hold for ImpactPlate and DeformingPlate (cf. Figure˜11 in Appendix˜E). Figure˜14 in Appendix˜E visualizes how high frequency errors accumulate over the rollout if we skip the later denosing steps, demonstrating the importance of reducing the high frequency errors for long rollouts despite low displacement RMSE.
Figure 6: Comparison of the predicted rollout displacements and displacement gradients on BendingBeam for different truncation steps $k_{\text{tr}}$. While the first diffusion steps strongly decrease the displacement RMSE, the later steps are important to reduce the local displacement gradient RMSE. A fast inference with a low denoising stride neither increases the displacement RMSE nor the gradient RMSE during a rollout.
[…]
@@
Ablations. As shown in Figure˜7, ablations confirm the importance of each ROBIN component. Reducing the diffusion length $K$ slightly decreases accuracy across all datasets. However, even with K=5K{=}5 diffusion steps, ROBIN remains significantly more accurate than all baselines. Using non-shared layers significantly degrades performance on BendingBeam, likely due to a reduced effective receptive field. The large error increase for State prediction indicates that residual prediction is crucial for Lagrangian simulations. Across all datasets, ROBIN outperforms the non-hierarchical variant, the non-diffusion variant, and the combination of DDPMs with the strongest hierarchical baseline (HCMT), demonstrating the synergy between DDPMs and AMPNs.
@@
Figure 7: Rollout RMSE of the displacement predictions. ROBIN remains accurate, even for a very small number of diffusion steps. Shared layers are crucial for BendingBeam to increase the receptive field. Replacing residual predictions with state predictions, hierarchical architectures with non-hierarchical architectures or with HCMT architectures, and diffusion with non-diffusion architectures significantly decrease the accuracy across all datasets.
Generalization to large meshes. Fine-tuning the pre-trained ROBIN converges markedly faster and attains substantially lower RMSE than training from scratch, demonstrating mesh-size independence and efficient transfer to deeper AMG hierarchies with twelve times more nodes, without architectural changes. In Figure˜8, pretrained predictions are visually indistinguishable from FEM even in thin, high-stress regions with large bending. The model trained from scratch exhibits negligible deformation due to much slower convergence within the same training budget.   ROBIN consistently requires $31$ s per case, while the numerical solver averages $108$ s and can require up to $4248$ s, due to problem-dependent nonlinear convergence costs. Table˜3 in Appendix˜E lists the quantitative results.
Figure 8: Comparison of the predicted rollout deformations and von Mises stresses (color, yellow is large) on BendingBeamLarge between a) the FEM, b) a fine-tuned ROBIN model, which was pre-trained on the small BendingBeam dataset, and, c)ROBIN model trained from scratch for the same number of training iterations. The fine-tuned ROBIN closely matches the FEM deformation and stress distribution, capturing even local high-stress hotspots, whereas the from-scratch model underestimates deformation due to much slower convergence within the same training budget.
@@ 6 Conclusion
Limitations and Future Work. ROBIN currently does not possess SO(3) equivariance. Adding this property could improve the accuracy of orientation-sensitive predictions. We focus on DDPM, while other diffusion formulations or denoising schedules could provide valuable insights and further enhance performance. Similarly, our experiments cover nonlinear solid mechanics, and extensions to other domains, such as fluid dynamics, are a promising direction. Lastly, ROBIN’s combination of fast inference and high accuracy opens opportunities for accelerating multi-stage design and optimization workflows.
@@ Acknowledgements
This work is part of the DFG AI Research Unit 5339 regarding the combination of physics-based simulation with AI-based methodologies for the fast maturation of manufacturing processes. The financial support by German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) is gratefully acknowledged. The authors acknowledge support by the state of Baden-Württemberg through bwHPC, as well as the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the German Federal Ministry of Education and Research. This work is supported by the Helmholtz Association Initiative and Networking Fund on the HAICORE@KIT partition.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The claims in the abstract and introduction are fully supported by our method section, as well as in the qualitative and quantitative results in the experiments section. The appendix provides more detailed results where required.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The conclusion discusses current limitations of the approach, including the scope of the paper and assumptions made.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not introduce any new theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We detail our method in Section 3, and provide additional information in the appendix where required. We detail our experimental setup and datasets used in the experiments section.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: We do not provide data and code at the time of submission, but will open-source both after acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify high-level training and test details in Section 4, and provide further details in Appendices DB and C.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report mean and standard deviation across five seeds for all main experiments. When evaluating the runtime-error tradeoff, we directly report information for all seeds without aggregation, explicitly showing the performance difference between runs.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide information on the compute resources for all experiments in Appendix C.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We have read the NeurIPS Code of Ethics. We made sure that our research complies to the Code of Ethics in every respect.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We include a discussion on broader impact in Appendix A.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We do not use pretrained language models, image generators or similar high-risk models in our approach, and do not scrape datasets. We still discuss potential cases for miss-use of our learned simulator in Appendix A.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We use two publicly available datasets, namely ImpactPlate and DeformingPlate. We credit the original papers of both explicitly.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: At time of submission, we do not release new assets. After acceptance, we will open-source our BendingBeam dataset, and provide appropriate documentation alongside it.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.